Do the following reactions favor reactants or products at equilibrium?



A. Sucrose (aq) + H2O(l) = Glucose (aq) + fructose (aq) k= 1.4 × 10^5



B. NH3 (aq) + H2O(l) = NH4^+(aq) + OH^-(aq) k= 1.6 × 10^-5



C. Fe2^03(s) + 3 co(g) = 2 Fe(s) + 3 co2(g) k( at 727°C)=24.2

1
Expert's answer
2021-12-15T10:04:17-0500

If the value of K is greater than 1, the products in the reaction are favored. If the value of K is less than 1, the reactants in the reaction are favored


A. Sucrose (aq) + H2O(l) = Glucose (aq) + fructose (aq) k= 1.4 × 10^5

Products are favoured


B. NH3 (aq) + H2O(l) = NH4^+(aq) + OH^-(aq) k= 1.6 × 10^-5

Reactants are favoured


C. Fe2^03(s) + 3 co(g) = 2 Fe(s) + 3 co2(g) k( at 727°C)=24.2

Products are favoured